OpenStudy (nottim):

how do you prove?: (1+cscA)/secA = cosA + cotA

Nnesha (nnesha):

what's the definition of csc and sec ? rewrite csc and sec in terms of sin cos.

OpenStudy (nottim):

(1/(1/sin))/(1/cos) = cos + cos/sin

Nnesha (nnesha):

when we have to verify the identity (prove) better way to do it is working on only one side either left or right our job is to prove L.H.S=R.H.S

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{ 1+\csc z }{\sec z} =\cos z+\cot z }\) that's your problem, And the simplest way is to recognize that; \(\color{#000000 }{ \displaystyle \cos z=\frac{1}{\sec z} }\) \(\color{#000000 }{ \displaystyle \cot z=\frac{\cos z}{\sin z}=\cos z\csc z =\frac{\csc z}{\sec z} }\)

Nnesha (nnesha):

so i suggest you to solve left side to prove its equal to right

Nnesha (nnesha):

and don't forget the plus sign `1+cscA` so \[\large\rm \color{Red}{\frac{1+cscA}{secA}} =cosA+cotA\] \[\large\rm \color{Red}{\frac{1+\frac{1}{sinA}}{\frac{1}{cosA}}} =cosA+cotA\] now just simplify the fraction

Nnesha (nnesha):

make sense ?

OpenStudy (nottim):

yes, thank you

Nnesha (nnesha):

did you get the same thing at left sidE ? both sides are equal ?